# Meta-Generalized-Gradient Approximation made Magnetic

### v2

@@ Abstract
where $\tau_{\rm unif}=\frac{3}{10}(3\pi^{2})^{2/3}\,n^{5/3}$ is the kinetic energy density of the uniform gas at density $n$. In particular, the SCAN form is defined by setting $F_{\mathrm{xc}}^{\mathrm{SCAN}}(n,s,\alpha)=h_{\mathrm{xc}}^{1}(n,s)+\left[h_{\mathrm{xc}}^{0}(n,s)-h_{\mathrm{xc}}^{1}(n,s)\right]f_{\mathrm{xc}}(\alpha)$ where $f_{\mathrm{xc}}(\alpha)$ interpolates between the semi-local xc enhancement factors for single-orbital densities ($h_{\mathrm{xc}}^{0}$) and for slowly-varying densities ($h_{\mathrm{xc}}^{1}$). This modeling exploits the fact that $\alpha$ [see Eq. (6) ] tends to zero in the iso-orbital limit (i.e., when the density is dominated by a single occupied orbital), tends to one in the uniform-density limit, and becomes $\gg 1$ in regions of low density between closed shells.
@@
in which $\tau_{\rm unif}^{\pm}=3/10(3\pi^{2})^{2/3}(n\pm|\vec{m}|)^{5/3}$, and all uniform gas quantities have been upgraded to an nc form by locally quantizing spin along the axis defined by the spin density $\vec{m}$.Kübler et al. (1988); not (d) The derivation of Eqs. (14) and (15) is provided in the Supplementary Material (this includes Ref.s Giuliani and Vignale, 2005; Lehtola et al., 2020; Desmarais et al., 2021; Vilela Oliveira et al., 2019; Peterson et al., 2007; Dolg, 2005; Balabanov and Peterson, 2005; Wilson et al., 1996; Campetella et al., 2020; Saunders et al., 1992; Towler et al., 1996; Doll et al., 2001, 2006; Desmarais et al., 2023; Yang et al., 2018; Wenzel et al., 2022; El Idrissi et al., 1991; Jones et al., 2022; Huber, 1979).ESI ; not (e)
@@
$\widetilde{\alpha}:=\frac{n^{\uparrow}\tau^{\uparrow}+n^{\downarrow}\tau^{\downarrow}-\frac{\left(\nabla n^{\uparrow}\right)\cdot\left(\nabla n^{\downarrow}\right)}{4}}{\frac{n^{\uparrow}}{2}\tau^{\uparrow}_{\rm unif}+\frac{n^{\downarrow}}{2}\tau^{\downarrow}_{\rm unif}}\;,$ (15)
where $\tau^{\sigma}_{\rm unif}=3/5(6\pi^{2})^{2/3}(n^{\sigma})^{5/3}$.not (f) Finally, the mSCAN reduces to the SCAN on closed-shell solutions (more below).

### v3

@@ Abstract
where $\tau_{\rm unif}=\frac{3}{10}(3\pi^{2})^{2/3}\,n^{5/3}$ is the kinetic energy density of the uniform gas at density $n$. In particular, the SCAN form is defined by setting $F_{\mathrm{xc}}^{\mathrm{SCAN}}(n,s,\alpha)=h_{\mathrm{xc}}^{1}(n,s)+\left[h_{\mathrm{xc}}^{0}(n,s)-h_{\mathrm{xc}}^{1}(n,s)\right]f_{\mathrm{xc}}(\alpha)$ where $f_{\mathrm{xc}}(\alpha)$ interpolates between the semi-local xc enhancement factors for single-orbital densities ($h_{\mathrm{xc}}^{0}$) and for slowly-varying densities ($h_{\mathrm{xc}}^{1}$). This modeling exploits the fact that $\alpha$ [see Eq. (6) ] tends to zero in the iso-orbital limit (i.e., when the density is dominated by a single occupied orbital), tends to one in the uniform-density limit, and becomes $\gg 1$ in regions of low density between closed shells
@@
in which $\tau_{\rm unif}^{\pm}\equiv 3/10(3\pi^{2})^{2/3}(n\pm|\vec{m}|)^{5/3}$, and all uniform gas quantities have been upgraded to an nc form by locally quantizing spin along the axis defined by the spin density $\vec{m}$.Kübler et al. (1988); not (d) The derivation of Eqs. (14) and (15) is provided in the Supplementary Material (this includes Ref.s Giuliani and Vignale, 2005; Oliver and Perdew, 1979; Lehtola et al., 2020; Desmarais et al., 2021; Vilela Oliveira et al., 2019; Peterson et al., 2007; Dolg, 2005; Balabanov and Peterson, 2005; Wilson et al., 1996; Campetella et al., 2020; Saunders et al., 1992; Towler et al., 1996; Doll et al., 2001, 2006; Desmarais et al., 2023; Dachsel et al., 1999; Yang et al., 2018; Wenzel et al., 2022; El Idrissi et al., 1991; Jones et al., 2022; Huber, 1979).ESI ; not (e)
@@
$\widetilde{\alpha}\equiv\frac{n^{\uparrow}\tau^{\uparrow}+n^{\downarrow}\tau^{\downarrow}-\frac{\left(\nabla n^{\uparrow}\right)\cdot\left(\nabla n^{\downarrow}\right)}{4}}{n^{\uparrow}\tau^{\uparrow}_{\rm unif}+n^{\downarrow}\tau^{\downarrow}_{\rm unif}}\;,$ (15)
where $\tau^{\sigma}_{\rm unif}=3/10(6\pi^{2})^{2/3}(n^{\sigma})^{5/3}$.not (f) Finally, the mSCAN reduces to the SCAN on closed-shell solutions (more below).
@@ V Pd Cr2
A word of caution and of encouragement are in order: The novel functional behaves differently from SCAN not only for magnetized states but also for open-shell states, therefore atomization energies should be scrutinized with care. We have also pointed out that band gaps may be worsened. Therefore, superseding the SCAN functional in all respects will require further efforts, which, we believe, can exploit the recently discovered extra-degrees of freedom that are allowed by the MGGAs. Lebeda et al. (2024)